An 8-bit register contains 87h. If this computer represents the numbers in 2’s complement system, what...

  1. An 8-bit register contains 87h. If this computer represents the numbers in 2’s complement system, what negative number in decimal does these 8 bits represent?

Please show your work step by step
Answer #1

87H = 1000 0111

Clearly, this number is negative since the MSB bit is set to 1.

Now, in order to find the magnitude of the number, we would have to find the 2's complement of 1000 0111.

1's complement of 1000 0111 = 0111 1000

2's complement  of 1000 0111 = 0111 1000 + 1 = 0111 1001

Now, we would need to find the decimal equivalent of 0111 1001.

0111 1001 = 0*27 + 1*26 + 1*25 + 1*24 + 1*23 + 0*22 + 0*21 + 1*20

= 64 + 32 + 16 + 8 + 1 = 121

Hence, the decimal number is = -121